OpenStudy (anonymous):

The triangles shown are similar. What is the value of x? A. 15 B. 12 C. 10 D. 9

OpenStudy (automaticloveletter):

Where are the triangles?

OpenStudy (automaticloveletter):

C..?

OpenStudy (automaticloveletter):

Yeah I'd go with C. First triangle has 4 as a side and the other one has 8 (4*2=8) First triangle has 3 as a side and the other one has 6 (3*2=6) Then the triangle has 5 as a side and the other one has 10 (5*2=10) If I'm wrong I'm sorry.